OpenStudy (anonymous):

how do you explain how two different equations have real rational and real irrational solutions? the equations are f(x) = x^2+6x-16 and g(x) = x^2 + 6x + 1

OpenStudy (jhannybean):

discriminant: \(d= b^2-4ac \) \(d > 0\) : 2 real and unreal roots \(d =0\) : only 1 real root. \(d <0\) : pair of complex roots in the form \(a\pm bi\) where \(i^2=-1\)

OpenStudy (jhannybean):

oh and \(d >0\) for perfect squares means you will have 2 rational roots.

OpenStudy (jhannybean):

Both of your equations are in quadratic form: \(ax^2+bx+c=0\), thus you have a discriminant: \(b^2-4ac\). You can use your discriminant to find your roots, to see whether they are complex or rational.

OpenStudy (anonymous):

thank you so much !